# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ Abstract
We study how a decision maker (DM) learns about the biases of unfamiliar information sources. Absent any friction, a rational DM uses familiar sources as yardsticks to discern the true bias of a source. If the DM has misspecified beliefs, this process fails. We derive long-run beliefs, behavior, welfare, and corresponding comparative statics when the DM holds dogmatic, incorrect beliefs about the biases of several familiar sources. The distortion due to misspecified learning aggregates misspecifications associated with familiar sources independently of sources the DM learns about. This has implications for labor market discrimination, media bias, and project finance and oversight.
@@ 1 Introduction
Good decision making in today’s information landscape involves the utilization of often biased information from a variety of sources. Some sources, such as traditional data or trusted advisors, are familiar to the decision maker (DM), and so he believes that he understands their biases well. Still others, such as algorithm recommendations and robo-advisors, are novel and unfamiliar, and the DM is much less certain about their biases. In an ideal setting, an astute decision maker might be able to learn about the bias inherent in each information source over time, using familiar sources to gauge unfamiliar ones, and his beliefs would eventually converge to the truth. However, a serious problem arises when the DM holds dogmatic, wrong, beliefs about the biases inherent in familiar sources, due to factors that elude detection and corrections. The long-run implications of this error are the subject of our paper.
@@
Following the paradigm of misspecified Bayesian learning pioneered by Berk (1966), we model a DM who consults several information sources to predict an evolving state of the world. Initially, the DM holds different prior beliefs about the biases of familiar and unfamiliar sources. For familiar sources, the prior is precise but can be inaccurate (hereafter misspecified), in that it concentrates on an event in which the perceived bias of the source differs from the truth. For unfamiliar sources, the prior is uncertain and, crucially, contains truth in its support. Over time, the DM observes signal realizations and updates his belief about source biases in a Bayesian manner. We examine how learning in the presence of multiple, misspecified sources breeds persistent misperceptions about unfamiliar sources and fuels distortions of the DM’s long-run behavior.
@@
Our framework can potentially be applied to a broad set of important economic phenomena. To illustrate, consider an upper-level manager who inquires about the productivity of minority employees from a mix of familiar sources — his personal source, senior raters, standard HR algorithms — and unfamiliar sources — novel employee data and newly installed committee members. Although these sources are prone to gender and racial biases, in a frictionless world, biases may be inconsequential, at least in the long run. For example, if a novel source gives consistently lower employee ratings than a known, impartial source, then it is regarded as adversarial to minorities, and its bias is accounted for when information from diverse sources is aggregated to predict minority employees’ productivity.
Reality, however, is not frictionless, particularly when the manager’s beliefs about familiar sources are misspecified. As someone who exhibits an implicit bias against minorities, the manager may give unduly low ratings to minority employees automatically and unconsciously. [^4] Implicit bias is an unconscious negative attitude towards a particular social group, making it challenging to eliminate due to its automatic nature. The renowned Implicit Association Test (IAT) developed by Greenwald et al. (1998) quickly assesses individuals’ associations between social groups and stereotypes, often in milliseconds. The opinions he gleans from other evaluators can be tainted by prejudice, stereotypes, and intergroup preferences, while his success in detecting these biases is limited, particularly in large organizations or when transparency and accountability are lacking (Roberson et al., 2007). Finally, algorithmic biases may arise due to poor designs or incomplete training, but they remained unnoticed until they garnered recent attention from academia and the popular press (Dastin, 2022). Since all these errors influence the manager’s learning about unfamiliar sources and, ultimately, the perceived productivity of minority employees and their compensation, their theoretical and practical consequences warrant a comprehensive investigation.
@@
Multi-sourced misspecified learning could be challenging to study for several reasons. In our leading example, suppose that the manager perceives two sources as impartial, although in reality one is biased in favor of minorities and the other is biased against them. Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences. One potential approach is to find a middle ground and use this as a neutral point to form beliefs about other sources. Still, the question remains: How exactly can this be achieved and how does the answer depend on the correlations between the signals? Correlation affects how different signals are weighed in the learning process. With multi-sourced misspecifications, the need to weigh different sources promotes the role of correlation in shaping learning outcomes, and the compound effect can escape even the most sophisticated analysis.
We develop an approach that works for arbitrary numbers of misspecified sources. Our analysis is conducted in a Gaussian environment in which the signal generated by a source in a given period equals the bias of the source, plus the true state of the world in that period and a Gaussian error. Errors are independent of the true state and of each other in the baseline model, but can take more general forms in later sections. Our main result, presented in Section 3.2, captures the distortions of the DM’s long-run perceptions and behavior by a single, intuitive, metric. In the baseline model, this metric equals the total misspecification associated with familiar sources, weighted by their relative precision to each other. As we exacerbate the misspecification associated with a source or increase its precision relative to others, its impact on the metric increases. In the leading example, if a majority of familiar sources, especially those exhibiting high signal-to-noise ratios, are biased against minorities but are nevertheless perceived as impartial, then the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities and is mirrored in his beliefs about unfamiliar sources through misspecified learning. In the long run, the manager underestimates the bias of any adversarial source toward minorities while overestimating that of any source favorable to them. He also underestimates minority employees’ productivity and pays them less than under the correctly specified model by our metric.
Our approach utilizes existing mathematical results on low-rank updates of invertible matrices. In the baseline model, signals (net of their biases) equal the true state plus idiosyncratic errors, so their covariance matrix is a rank-one update of a diagonal matrix. When signals are correlated through a single common channel, the impact of correlation on the outcomes of misspecified learning is limited. Consequently, the metric that captures the distortions of the DM’s long-run belief and behavior exhibits a striking feature that we call DOOM, as it depends only on sources with misspecified biases but not those that the DM learns about. DOOM continues to hold in variants of the baseline model in which the signals depend on a common latent factor, or when the true state and the signals can take high-dimensional values (see Sections 4.1 and 4.2). When correlations between signals become arbitrary, the DM’s misperceptions about source biases satisfy a weaker form of DOOM, as they depend only on misspecified sources and their correlations with unfamiliar sources, but nothing else. Further, the distortion in the DM’s long-run behavior always satisfies DOOM, due to the simple fact that the DM’s limiting belief concentrates on the minimizer of the Kullback-Leibler divergence from truth.
@@
The comparative statics of our metric have important implications for reality. In the context of labor market discrimination, DOOM implies that the damage caused by a few rotten apples cannot be rectified by installing objectively neutral but lesser-known members of the minority group to the evaluation panel, or by training junior raters to reduce errors in employee assessments. The same cannot be said about sources that the DM holds dogmatic, wrong, beliefs about, as our metric is in general nonmonotonic in their signal-to-noise ratios. In the presence of multiple misspecified sources, improving the precision of one source renders the others relatively imprecise. The overall impact on our metric is in general ambiguous, and this has implications for the use of rater training to curtail discrimination. Details are in Section 3.5.
@@
While it is tempting to conjecture that correcting the misspecifications associated with familiar sources may improve learning outcomes, reality is more subtle when the covariance matrix of the signals is also subject to misspecified learning. In Section 4.3, we show that learning the covariance matrix of the signals, in addition to their biases, makes the signals appear noisier to the DM than they truly are. Consequently, the DM relies less on signal realizations when predicting the true state of the world, and this countervailing force is shown to partially correct the distortions due to misspecified learning. Policies that operate by correcting misspecifications block this novel channel and so could entail an ambiguous consequence on learning outcomes. This finding speaks to the debiasing programs used by real-world organizations to address discrimination, such as implicit bias training, priming shared identities, and promoting algorithmic fairness. It suggests that even if these programs are effective, their behavioral and welfare consequences may remain ambiguous.
In addition to employee evaluation and compensation, the aforementioned phenomenon occurs when voters form perceptions of the biases inherent in new social media sites based on their erroneous beliefs about traditional media. There is evidence that partisans tend to view party-congenial sources as neutral, while displaying excessive hostility toward out-party sources (Pennycook and Rand, 2021; Hassell et al., 2020). Likewise, a manager seeking project advice may allow his hardwired misperceptions about familiar sources — such as overoptimism, confirmation bias, and overtrust in one’s inner circle (Malmendier, 2018) — to shape his views about unfamiliar sources such as external consultants and robo-advisors. In Section 5, we demonstrate how these ingrained biases can spill into new domains, generating the hostile media effect and false polarization, [^5] The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is. False polarization measures the difference between the perceived distance between two groups and the actual distance between them. Both phenomena have attracted considerable scholarly attention within the domains of political science, communication, and psychology. Section 5.1 offers a synthesis of the literature. in addition to polarizing voters’ behavior and distorting managers’ portfolio choices. We further contend, based on our comparative statics, that corrective measures that target political misinformation or managerial biases may give rise to unintended consequences. For example, while fact-checking new media always improves voters’ welfare because of DOOM, fact-checking tradition media generally has an ambiguous impact due to the distortions that arise from misspecified learning. The latter are nonmonotonic in the accuracy of familiar sources and the treatment effects of misperception recalibration.
@@ 1.1 Related literature
The study of misspecified Bayesian learning models has a long tradition in statistical decision theory. The canonical paradigm endows the DM with a prior belief over the models that describe the underlying environment. If the prior does not contain the true model in its support, it is misspecified. The seminal work of Berk (1966) provides conditions under which the DM’s limiting posterior belief concentrates on the set of models that minimize the Kullback-Leibler (KL) divergence from truth.
Economic models of misspecified Bayesian learning can be broadly categorized into passive learning models and active learning models. The former, including ours, assume that the actions taken by the DM have no impact on the signals generated by truth. The work most related to ours is carried out by Heidhues et al. (2019) (HKS). These authors consider environments where the signals generated by unknown fundamentals follow multivariate Guassian distributions. Such environments are shown to satisfy Berk’s conditions for the convergence of the limiting belief to KL minimizers. Our proof of the convergence of the limiting belief builds on this result.
The main differences between our results and those of HKS lie in the characterization of the limiting belief. In particular, HKS focus on situations where a single dimension of the fundamentals is misspecified, which is the natural setting to understand the implications of overconfidence. We entertain the possibility of multidimensional misspecifications, which is essential to understand how a DM aggregates biased sources in the settings discussed previously. Our characterization extends beyond that of HKS and works for arbitrary numbers of misspecified sources with general covariance matrices. The structural insights it uncovers, such as DOOM, have escaped the attention of previous work. Our comparative statics are sharpest when the covariance matrix of the signals constitutes low-rank updates of diagonal matrices. HKS consider alternative forms of covariance matrices that are suitable for their exploration of the consequences of overconfidence on intergroup prejudice.
@@
In contrast, models of misspecified active learning allow the DM’s actions to affect the signals that he subsequently uses to form beliefs about the truth. The literature on misspecified active learning has recently advanced significantly. Among others, Esponda and Pouzo (2016) propose Berk-Nash equilibrium as a steady-state notion for misspecified learning dynamics. Esponda et al. (2021), Fudenberg et al. (2021), Bohren and Hauser (2021), and Frick et al. (2023) examine the convergence of Berk-Nash equilibria in single and multi-agent environments. Ba (2021) entertains the possibility of regime shifts. He and Libgober (2020) examine the evolutionary foundations of perpetual model misspecification. None of these papers shares our focus on the characterization of steady states with multi-sourced misspecifications.
There is a broader conceptual connection between our work and the Bayesian network approach of Spiegler (2020) to model misspecified causality, in that both allow for misspecifications of high-dimensional objects. Yet Spiegler’s use of directed acyclic graphs to model causality has no obvious counterpart in our setting, and his solution concept requires that the DM take endogenous actions to fully explain the observed data. Our steady state does not involve endogenous actions and has a misspecified-Bayesian-learning foundation.
Developments on the theoretical frontier have spurred novel research in various areas, including labor market discrimination, media bias, and finance. Within the realm of labor market discrimination, Bohren et al. (2019) and Bohren et al. (2023) propose methods to distinguish discrimination due to erroneous beliefs from those rooted in tastes or rational beliefs. The current paper adds to this agenda by exploring a novel channel: misspecified learning of unfamiliar information sources. The main takeaways from our analysis — namely dogmatic, wrong, beliefs about familiar sources can breed persistent misperceptions about unfamiliar sources and generate nuanced comparative statics when subject to informational interventions — have no counterparts in existing theories of discrimination. To the extent that interventions such as improved source accuracy or misperception recalibration are also proposed by Bohren et al. (2023) to identify and address discrimination, our analysis could impact how policies are designed to combat discrimination in the real world.
The economic literature on media bias is vast, but is devoted to understanding actual media biases (Anderson et al., eds, 2016).666(Mis)perceived media bias is an important topic in political science, communication, and related fields (Feldman, 2014). These strands of literature are discussed in Section 5.1. A notable exception is made by Gentzkow et al. (2018) (GWZ), whose model of misperceived media bias features a payoff-irrelevant confounder that is independent of the payoff-relevant state. A source is biased if it exhibits a nonzero correlation with the confounder. The DM misperceives his personal source as unbiased and uses this belief to guide his consumption of other media sources.
The model of GWZ differs from ours in its construct and predictions, in addition to requiring different methodologies to analyze. First, GWZ examine an active learning model with a single misspecified source. Second, their way of modeling biases (as correlations with a random confounder) implies that the distortion in the DM’s long-run behavior has a zero unconditional mean, which is not the case here. A central focus of GWZ is on the endogenous trust in news sources, defined as the latter’s (mis)perceived correlations with the true state. Trust is correctly specified throughout this paper and is known to the DM for the most part. When endogenized as a subject of misspecified learning in Section 4.3, we find no effect on perceived news biases and only a moderation of the DM’s behavior.
The literature on behavioral finance is vast (see Malmendier 2018 for a survey), although research that explores misspecified Bayesian learning as a source of friction remains sparse. Molavi et al. (2023) examine a passive learning model in which asset returns follow a high-dimensional hidden factor model, although the models used by market participants to explain observed data have a lower dimensionality. The authors examine the consequences of this departure from the rational expectations paradigm for asset pricing and the dynamics of forecast errors. The Sherman-Morrison formula — a tool to invert rank-one perturbations of invertible matrices — is present in their mathematical derivations (and those of Levy et al. 2022) but plays a less significant role than in our analysis.
@@ 2 Baseline model
Time is discrete and infinite. In each period $t=1,2,\cdots$, a random state $\omega_{t}$ is drawn independently from the standard normal distribution. A decision maker (DM) — informed by a finite set $\mathcal{S}\coloneqq\{1,\cdots,N\}$ of sources — predicts the mean of $\omega_{t}$. The signal generated by source $i\in\mathcal{S}$ in period $t$ is $X_{i,t}=\omega_{t}+b_{i}+\varepsilon_{i,t}$, where $b_{i}\in\mathbb{R}$ is the bias of the source, and $\varepsilon_{i,t}$ is a Gaussian error with mean zero and variance $v_{i}>0$. In the baseline model, errors are assumed to be independent of the true state, across each other, and over time.
@@
Initially, the DM holds a prior belief about source biases. In every subsequent period, he updates his belief in a Bayesian manner, based on the prior and signal realizations. It is well known that if true biases, which we denote by $b_{i}^{*}$s, lie in the support of the prior, then the DM’s posterior belief will eventually converge to the truth. His long-run behavior is the best linear predictor of the state given signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$, under the (correct) belief that source biases have value $b^{*}\coloneqq[b_{1}^{*},\text{ }\cdots\text{ },b_{N}^{*}]^{\top}$.
@@
We examine the case where the DM’s prior is partly misspecified. In the baseline model, we focus on a particular kind of misspecification, whereby the DM holds dogmatic, wrong, beliefs about the biases of some sources and uses them to guide his learning about other sources. We examine how such misspecifications distort the DM’s learning and, ultimately, his long-run belief and behavior.
@@
Formally, we say that the DM misspecifies the bias of source $i$ if he initially assigns probability one to an event in which $b_{i}=\tilde{b}_{i}$, for some $\tilde{b}_{i}\neq b_{i}^{*}$. Let $\mathcal{M}\coloneqq\{1,\cdots,M\}$ denote the set of sources with misspecified biases. Suppose that $1\leq M<N$, so that the DM misspecifies the biases of some but not all sources. For those sources in $\mathcal{S}-\mathcal{M}\coloneqq\{M+1,\cdots,N\}$, we assume that the DM’s prior is nondegenerate on $\mathbb{R}^{N-M}$. The support of the prior is $\Theta\coloneqq\{b\in\mathbb{R}^{N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$. The outcomes of Bayesian learning based on this prior are the subject of the current study.
@@
Three remarks are in order. For now, we assume that signals equal biases plus the true state and independent errors, and that their covariance matrix is known to the DM from the outset. In Sections 3.4 and 4.3, we entertain more general forms of signals whose covariance matrix may also be a subject of misspecified learning. Second, we treat misspecification and degenerate, wrong, prior as synonyms for most of the paper; alternative forms of misspecifications are discussed in Section 4.1. Finally, it will become clear that adding sources with correctly specified biases to the analysis would not affect our predictions in any meaningful way.

### v3

# Learning Source Biases: Multisource Misspecifications and Their Impact on Predictions

@@ Abstract
We study how a Bayesian decision maker (DM) learns about the biases of novel information sources to predict a random state. Absent frictions, the DM uses familiar sources as yardsticks to accurately discern the biases of novel sources. We derive the distortion of the DM’s long-run prediction when the DM holds misspecified beliefs about the biases of several familiar sources. The distortion aggregates misspecifications across familiar sources independently of the number and nature of the novel sources the DM learns about. This has implications for labor market discrimination, media bias, and project finance and oversight.
@@ 1 Introduction
Making good predictions in today’s information landscape requires a decision maker (DM) to synthesize information from a variety of biased sources. Some sources, such as traditional data or trusted advisors, are familiar to the DM, and so he believes that he understands their biases well. Still others, such as algorithm recommendations and robo-advisors, are novel and unfamiliar, and the DM is much less certain about their biases. In an ideal setting, an astute DM might be able to learn about the bias inherent in each information source over time — using familiar sources to gauge unfamiliar ones — and his long-run prediction will be accurate. However, a serious problem arises when the DM holds dogmatic, wrong, beliefs about the biases inherent in familiar sources, due to factors that elude detection and corrections. The long-run implications of this error are the subject of our paper.
@@
Following the paradigm of misspecified Bayesian learning, we model a DM who consults several information sources to predict an evolving state of the world. Initially, the DM holds different prior beliefs about the biases of familiar and unfamiliar sources. For familiar sources, the prior is precise but can be inaccurate (hereinafter misspecified), in that it concentrates on an event in which the perceived bias of the source differs from the truth. For unfamiliar sources, the prior is uncertain and, crucially, contains the truth in its support. Over time, the DM observes signal realizations and updates his beliefs about source biases in a Bayesian manner. We examine how learning in the presence of multiple, misspecified sources breeds persistent misperceptions about unfamiliar sources and distorts the DM’s long-run prediction.
@@
Our framework can be applied to a broad set of important economic phenomena. To illustrate, consider an upper-level manager who inquires about the productivity of minority employees from a mix of familiar sources — his personal source, senior raters, standard HR algorithms — and unfamiliar sources — novel employee data and newly installed committee members. Although these sources are prone to gender and racial biases, such biases may be inconsequential in the long run absent learning frictions: if a novel source consistently gives lower ratings to minorities than a known, impartial source, then its negative bias can be accounted for when information is aggregated to predict minority employees’ productivity. In reality, however, the manager may hold misspecified beliefs about familiar sources, due to his implicit bias against minorities (Greenwald et al., 1998), or his ignorance of the prejudice, stereotypes, and intergroup preferences inherent in the evaluations of other raters or algorithms (Roberson et al., 2007; Dastin, 2022). Since these unconscious biases can result in persistent underprediction and underpayment of minority employees, their theoretical and practical implications warrant a comprehensive investigation.
@@
Our analysis is conducted in a Gaussian environment in which the signal generated by a source in a given period equals the bias of the source, plus the true state of the world in that period and an error. Errors are independent of the true state but can exhibit arbitrary correlations among themselves. Studying multi-sourced misspecified learning in this environment could be challenging. In our leading example, suppose that the manager perceives two sources as impartial, although in reality one is biased in favor of minorities and the other is biased against them. Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences. One potential approach is to aggregate the differences between these sources and use the outcome as a middle ground to form beliefs about other sources and the true state. Still, the question remains: How exactly can this be achieved in general environments where familiar sources, unfamiliar sources, and the true state may exhibit arbitrary correlations?
Our main result, presented in Section 3.2, describes how the misspecifications associated with familiar sources are aggregated into distortions of the DM’s long-run beliefs about source biases and his prediction of the true state. Strikingly, the distortion in his prediction does not depend on unfamiliar sources at all, but instead depends only on misspecified sources, a property we call DOOM. The proof of DOOM exploits fundamental properties of the statistical decision problem faced by the DM, which has two stages. In the first stage, the DM learns the biases of unfamiliar sources using misspecified, familiar sources as yardsticks, which amounts to minimizing the Kullback-Leibler (KL) divergence from the truth. In linear-Gaussian environments, this is equivalent to forming the constrained least squares estimator of the biases of unfamiliar sources (at population level). In the second stage, the DM uses the first-stage result to form the best linear predictor of the true state. Since the former is the best constrained linear estimator of source biases, distortions thereof — which have already been optimized — should not further distort the second-stage predictor, regardless of the number and nature of the unfamiliar sources being learned.
We can draw sharper results by imposing additional structures on the covariance matrix. In Section 3.3, we examine a simple case in which the errors associated with the signals are jointly independent. Leveraging matrix algebra theories on rank-one updates of matrices, we develop a single metric that fully captures the distortionary effects of misspecified learning. The metric equals the total misspecification associated with familiar sources, weighted by their relative precision. In our leading example, it takes the form an unconscious negative bias against minorities and is mirrored in the manager’s beliefs about unfamiliar sources through misspecified learning. In the long run, the DM underestimates the bias of any unfamiliar source that is adversarial toward minorities while overestimating the bias of any unfamiliar source that is favorable to them. Consequently, he underpredicts minority employees’ productivity and pays them less than under the correctly specified model.
@@
The comparative statics of our distortionary metric have important implications for reality. In the context of labor market discrimination, DOOM implies that the damage caused by a few rotten apples cannot be rectified by installing objectively neutral but lesser-known members of the minority group to the evaluation panel, or by training junior raters to reduce rating errors. The same cannot be said about sources that the DM holds dogmatic, wrong, beliefs about, as our metric is nonmonotonic in their signal-to-noise ratios in general. In the presence of multiple misspecified sources, improving the precision of one source renders the others relatively imprecise. The overall impact on our metric is generally ambiguous.
@@
While it is tempting to conjecture that correcting the misspecifications associated with familiar sources may improve predictions, reality is more subtle when the covariance matrix of the signals is also subject to misspecified learning. In Section 5.1, we show that learning the covariance matrix of the signals, in addition to their biases, makes the signals appear noisier to the DM without further distorting his beliefs about unfamiliar sources compared to the baseline model. Consequently, the DM relies less on signal realizations than before when predicting the true state, and this countervailing force is shown to partially mitigate the distortions caused by misspecified learning. Policies aimed at correcting misspecifications may undermine this bias mitigation effect, resulting in an exacerbation of distortions. This finding speaks to the debiasing programs used by real-world organizations to curb implicit discrimination (e.g., implicit bias training, efforts to reduce intergroup animosity and to promote algorithmic fairness), suggesting that even if these programs are effective, their behavioral and welfare consequences may remain ambiguous.
Our model can additionally be applied to study misperceived media bias and project management under misspecified learning. We summarize our findings — particularly the implications of DOOM for polarization, fact-checking and the efficacy of corporate repair policies — in Section 6 and defer the analysis to the Online Appendix. The desirability of DOOM, including its use to identify misspecified sources, its relevance to alternative decision situations, and its robustness to alternative forms of misspecifications and high-dimensional states and signals, are explored in Sections 4 and 5.2.
@@ 1.1 Related literature
The study of misspecified Bayesian learning models has a long tradition in statistical decision theory. The seminal work of Berk (1966) provides conditions under which the limiting posterior beliefs concentrate on the model that minimizes the Kullback-Leibler divergence from the truth. Economic models of misspecified Bayesian learning can be broadly categorized into passive and active learning models. The former, including ours, assume that the actions taken by the DM do not affect the signals generated by the underlying truth. Our contributions to passive misspecified learning are twofold. First, we propose a model of making predictions using familiar and unfamiliar information sources applicable to broad phenomena. Second, we obtain general characterizations of the long-run outcomes using fundamental principles from statistical decision making.
The most related work to ours is Heidhues et al. (2019) (hereinafter, HKS), which examines a passive learning model with a single source of misspecification and demonstrates that Berk’s convergence result applies in multidimensional Gaussian environments. Our convergence proof builds on their analysis. The main difference between HKS and the current work lies in the DM’s decision problem. HKS examine how the DM’s overconfidence in his caliber distorts his estimates about ingroups and outgroups — a problem akin to what our DM faces in the first stage. However, our DM uses multiple misspecified sources to guide his learning of unfamiliar sources and, crucially, uses the learned model to predict a random state. This second stage, absent from HKS, generates new insights such as DOOM that have eluded previous researchers.
Applications of misspecified passive learning models to other areas of economics include Bohren et al. (2019) and Molavi et al. (2023). Bohren et al. (2019) examine, theoretically and experimentally, a model in which a DM holds erroneous prior beliefs about minorities and updates his beliefs over time based on new information. The authors characterize the trajectory of beliefs and distinguish their result from traditional taste-based or rational statistical discrimination. Our model addresses labor market discrimination in the long run and its relationship between traditional and novel information sources. Molavi et al. (2023) examine a model in which asset returns follow a high-dimensional hidden factor model, while market participants are restricted to using lower-dimensional models to explain observed data. Their misspecification differs from ours in nature, and they explore its consequences for asset pricing and the dynamics of forecast errors.
@@
Models of misspecified active learning allow the DM’s actions to affect the signals used to form beliefs about the truth. The theory of misspecified active learning has recently advanced significantly, with a primary focus on the convergence properties of various solution concepts. [^5] Among others, Esponda and Pouzo (2016) propose Berk-Nash equilibrium as a steady state notion for misspecified learning dynamics. Fudenberg et al. (2021) and Frick et al. (2023) examine the convergence of Berk-Nash equilibria in single and multi-agent environments. Ba (2021) entertains the possibility of regime shifts. He and Libgober (2020) examine the evolutionary foundation of perpetual model misspecification. There is also a broader conceptual connection between our work and the Bayesian network approach of Spiegler (2020) to model misspecified causality. Both allow for misspecifications of high-dimensional objects, but Spiegler’s use of directed acyclic graphs to model causality has no obvious counterpart in our setting, and his solution concept requires that the DM take endogenous actions to fully explain the observed data.
@@ 2 Baseline model
Time is discrete and infinite. In each period $t\in\mathbb{N}$, a random state $\omega_{t}$ is drawn independently from the standard normal distribution. A DM — informed by a finite set $\mathcal{S}\coloneqq\{1,\cdots,N\}$ of sources — predicts the mean of $\omega_{t}$. The signal generated by source $i\in\mathcal{S}$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\epsilon_{i,t}$, where $b_{i}\in\mathbb{R}$ is the bias of the source, and $\epsilon_{i,t}$ is a state-independent Gaussian error with mean zero and variance $v_{i}>0$.
@@
Initially, the DM holds prior beliefs about source biases. In every subsequent period, he updates his beliefs in a Bayesian manner based on the prior and signal realizations. It is well known that if true biases, which we denote by $b_{i}^{*}$, $i\in\mathcal{S}$, lie in the support of the prior, then the DM’s posterior beliefs will eventually concentrate on the truth. His long-run behavior is the best linear predictor of the state given signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$, under the correct belief that source biases have value $b^{*}\coloneqq[b_{1}^{*},\text{ }\cdots\text{ },b_{N}^{*}]^{\top}$.
@@
We examine the case where the DM’s prior is partly misspecified. In the baseline model, we focus on a particular kind of misspecification, whereby the DM holds dogmatic, wrong, beliefs about the biases of some sources and uses them to guide his learning about other sources. We examine how such misspecifications distort the DM’s learning and, ultimately, his long-run beliefs and prediction.
@@
Formally, we say that the DM misspecifies the bias of source $i$ if he initially assigns probability one to an event in which $b_{i}=\tilde{b}_{i}$, for some $\tilde{b}_{i}\neq b_{i}^{*}$. Let $\mathcal{M}\coloneqq\{1,\cdots,M\}$ denote the set of sources with misspecified biases. Suppose that $1\leq M<N$, so that the DM misspecifies the biases of some but not all sources. For those sources in $\mathcal{S}-\mathcal{M}\coloneqq\{M+1,\cdots,N\}$, we assume that the DM’s prior is nondegenerate on $\mathbb{R}^{N-M}$. The support of the prior is $\Theta\coloneqq\{b\in\mathbb{R}^{N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$. The long-run outcomes of learning based on this prior are the subject of the current study.
@@
Several remarks are in order. First, the assumption of unit covariance between the true state and the signals is made to simplify the exposition. The case of general, known, covariances is discussed in Section 3.2. Second, we treat misspecification and degenerate, wrong, prior as synonyms for most of the paper; alternative forms of misspecifications are discussed in Section 4. Third, we assume, for now, that the true covariance matrix $\Omega^{*}$ of the errors — which can be any symmetric and positive definite matrix in $\mathbb{R}^{N\times N}$ — is known to the DM. In Section 5.1, we consider the possibility that the covariance matrix may also be a subject of misspecified learning. Finally, it will become evident that adding sources with correctly specified biases to the analysis would not affect our predictions in any meaningful way.